OpenStudy (fwalk1ga39):

How would you combine like terms for -7+9n-8n

OpenStudy (fwalk1ga39):

And explain how

OpenStudy (haleyelizabeth2017):

Well, like terms would be for example \(5x\)+9\(+5x\) 5x and 5x would be the like terms. You'd just add them to get 10x+9 Think you have an idea of how to do yours?

rebeccaxhawaii (rebeccaxhawaii):

so 9n-8n=? then -7 as the constant

rebeccaxhawaii (rebeccaxhawaii):

example 4n-3n would be just n 4-3=1

OpenStudy (haleyelizabeth2017):

Hint: \[-7+\color {red}{9n-8n}\]

OpenStudy (fwalk1ga39):

I no the answer is n-7 I just have a hard time explaining the reasoning.

OpenStudy (haleyelizabeth2017):

Well if this is a step for something bigger, you could just say combine like terms. Just explain how 9-8 is 1 and just explain it that way

rebeccaxhawaii (rebeccaxhawaii):

bc 9n-8n=n or 1n just like 9-8=1

OpenStudy (fwalk1ga39):

They want you to explain every step and give reasoning.